# Node Duplication Improves Cold-start Link Prediction

### v1

@@ Abstract
Graph Neural Networks (GNNs) are prominent in graph machine learning and have shown state-of-the-art performance in Link Prediction (LP) tasks. Nonetheless, recent studies show that GNNs struggle to produce good results on low-degree nodes despite their overall strong performance. In practical applications of LP, like recommendation systems, improving performance on low-degree nodes is critical, as it amounts to tackling the cold-start problem of improving the experiences of users with few observed interactions. In this paper, we investigate improving GNNs’ LP performance on low-degree nodes while preserving their performance on high-degree nodes and propose a simple yet surprisingly effective augmentation technique called NodeDup. Specifically, NodeDup duplicates low-degree nodes and creates links between nodes and their own duplicates before following the standard supervised LP training scheme. By leveraging a “multi-view” perspective for low-degree nodes, NodeDup shows significant LP performance improvements on low-degree nodes without compromising any performance on high-degree nodes. Additionally, as a plug-and-play augmentation module, NodeDup can be easily applied on existing GNNs with very light computational cost. Extensive experiments show that NodeDup achieves 38.49%, 13.34%, and 6.76% improvements on isolated, low-degree, and warm nodes, respectively, on average across all datasets compared to GNNs and state-of-the-art cold-start methods.
Machine Learning, ICML
@@ 1 Introduction
Link prediction (LP) is a fundamental task of graph-structured data (Liben-Nowell & Kleinberg, 2007; Trouillon et al., 2016), which aims to predict the likelihood of the links existing between two nodes in the network. It has wide-ranging real-world applications across different domains, such as friend recommendations in social media (Sankar et al., 2021; Tang et al., 2022; Fan et al., 2022), product recommendations in e-commerce platforms (Ying et al., 2018; He et al., 2020), knowledge graph completion (Li et al., 2023; Vashishth et al., 2020; Zhang et al., 2020), and chemical interaction prediction (Stanfield et al., 2017; Kovács et al., 2019; Yang et al., 2021; Guo et al., 2022a).
@@
In recent years, graph neural networks (GNNs) (Kipf & Welling, 2016a; Veličković et al., 2017; Hamilton et al., 2017) have been widely applied to LP, and a series of cutting-edge models have been proposed (Zhang & Chen, 2018; Zhang et al., 2021; Zhu et al., 2021; Zhao et al., 2022b). Most GNNs follow a message-passing scheme (Gilmer et al., 2017) in which information is iteratively aggregated from neighbors and used to update node representations accordingly. Consequently, the success of GNNs usually heavily relies on having sufficient high-quality neighbors for each node (Zheng et al., 2021; Liu et al., 2021). However, real-world graphs often exhibit long-tailed distribution in terms of node degrees, where a significant fraction of nodes have very few neighbors (Tang et al., 2020b; Ding et al., 2021; Hao et al., 2021). For example, Figure 1 shows the long-tailed degree distribution of the Citeseer dataset. Moreover, LP performances w.r.t. node degrees on this dataset also clearly indicate that GNNs struggle to generate satisfactory results for nodes with low or zero degrees. For simplicity, in this paper, we refer to the nodes with low or zero degrees as cold nodes and the nodes with higher degrees as warm nodes.
To boost GNNs’ performance on cold nodes, recent studies have proposed various training strategies (Liu et al., 2020, 2021; Zheng et al., 2021; Hu et al., 2022) and augmentation strategies (Hu et al., 2022; Rong et al., 2019; Zhao et al., 2022b) to improve representation learning quality. For instance, ColdBrew (Zheng et al., 2021) posits that training a powerful MLP can rediscover missing neighbor information for cold nodes; TailGNN (Liu et al., 2021) utilizes a cold-node-specific module to accomplish the same objective. However, such advanced training strategies (e.g., ColdBrew and TailGNN) share a notable drawback: they are trained with a bias towards cold nodes, which then sacrifices performance on warm nodes (empirically validated in Table 1). However, in real-world applications, both cold nodes and warm nodes are critical (Clauset et al., 2009). On the other hand, while augmentation methods such as LAGNN (Liu et al., 2022b) do not have such bias, they primarily focus on improving the overall performance of GNNs in LP tasks, which may be dominated by warm nodes due to their higher connectivity. Additionally, the augmentation methods usually introduce a significant amount of extra computational costs (empirically validated in Figure 5). In light of the existing work discussed above on improving LP performance for cold nodes, we are naturally motivated to explore the following crucial but rather unexplored research question:
@@
We observe that cold node LP performance usually suffers because they are under-represented in standard supervised LP training due to their few (if any) connections. Given this observation, in this work, we introduce a simple yet effective augmentation method, NodeDup, for improving LP performance on cold nodes. Specifically, NodeDup duplicates cold nodes and establishes edges between each original cold node and its corresponding duplicate. Subsequently, we conduct standard supervised end-to-end training of GNNs on the augmented graph. To better understand why NodeDup is able to improve LP performance for cold nodes, we thoroughly analyze it from multiple perspectives, during which we discover that this simple technique effectively offers a “multi-view” perspective of cold nodes during training. This “multi-view” perspective of the cold nodes acts similarly to an ensemble and drives performance improvements for these nodes. Additionally, our straightforward augmentation method provides valuable supervised training signals for cold nodes and especially isolated nodes. Furthermore, we also introduce NodeDup(L), a lightweight variation of NodeDup that adds only self-loop edges into training edges for cold nodes. NodeDup(L) empirically offers up to a 1.3$\times$ speedup over NodeDup for the training process and achieves significant speedup over existing augmentation baselines. In our experiments, we comprehensively evaluate our method on seven benchmark datasets. Compared to GNNs and state-of-the-art cold-start methods, NodeDup achieves 38.49%, 13.34%, and 6.76% improvements on isolated, low-degree, and warm nodes, respectively, on average across all datasets. NodeDup also greatly outperforms augmentation baselines on cold nodes, with comparable warm node performance. Finally, as plug-and-play augmentation methods, our methods are versatile and effective with different LP encoders/decoders. They also achieve significant performance in a more realistic inductive setting.
@@ 2 Preliminaries
2.1 Notation
Let an attributed graph be $G=\{\mathcal{V},\mathcal{E},\textbf{X}\}$, where $\mathcal{V}$ is the set of $N$ nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edges where each $e_{vu}\in\mathcal{E}$ indicates nodes $v$ and $u$ are linked. Let $\textbf{X}\in\mathbb{R}^{N\times F}$ be the node attribute matrix, where $F$ is the attribute dimension. Let $\mathcal{N}_{v}$ be the set of neighbors of node $v$, i.e., $\mathcal{N}_{v}=\{u|e_{vu}\in\mathcal{E}\}$, and the degree of node $v$ is $|\mathcal{N}_{v}|$. We separate the set of nodes $\mathcal{V}$ into three disjoint sets $\mathcal{V}_{iso}$, $\mathcal{V}_{low}$, and $\mathcal{V}_{warm}$ by their degrees based on the threshold hyperparameter $\delta$ [^1] This threshold $\delta$ is set as 2 in our experiments, based on observed performance gaps in LP on various datasets, as shown in Figure 1 and Figure 6. Further reasons for this threshold are detailed in Section D.1.. For each node $v\in\mathcal{V}$, $v\in\mathcal{V}_{iso}$ if $|\mathcal{N}_{v}|=0$; $v\in\mathcal{V}_{low}$ if $0<|\mathcal{N}_{v}|\leq\delta$; $v\in\mathcal{V}_{warm}$ if $|\mathcal{N}_{v}|>\delta$. For ease of notation, we also use $\mathcal{V}_{cold}=\mathcal{V}_{iso}\cup\mathcal{V}_{low}$ to denote the cold nodes, which is the union of Isolated and Low-degree nodes.
2.2 LP with GNNs
In this work, we follow the commonly-used encoder-decoder framework for GNN-based LP (Kipf & Welling, 2016b; Berg et al., 2017; Schlichtkrull et al., 2018; Ying et al., 2018; Davidson et al., 2018; Zhu et al., 2021; Yun et al., 2021; Zhao et al., 2022b), where a GNN encoder learns the node representations and the decoder predicts the link existence probabilities given each pair of node representations. Most GNNs follow the message passing design (Gilmer et al., 2017) that iteratively aggregate each node’s neighbors’ information to update its embeddings. Without the loss of generality, for each node $v$, the $l$-th layer of a GNN can be defined as
$\bm{h}_{v}^{(l)}=\text{\small UPDATE}\big{(}\bm{h}_{v}^{(l-1)},\bm{m}_{v}^{(l-1)}\big{)},$
@@ 2.2 LP with GNNs
where $\bm{h}_{v}^{(l)}$ is the $l$-th layer’s output representation of node $v$, $\bm{h}_{v}^{(0)}=\bm{x_{v}}$, $\text{\small AGG}(\cdot)$ is the (typically permutation-invariant) aggregation function, and $\text{\small UPDATE}(\cdot)$ is the update function that combines node $v$’s neighbor embedding and its own embedding from the previous layer. For any node pair $v$ and $u$, the decoding process can be defined as $\hat{y}_{vu}=\sigma\big{(}\text{\small DECODER}(\bm{h}_{v},\bm{h}_{u})\big{)},$ where $\bm{h}_{v}$ is the GNN’s output representation for node $v$ and $\sigma$ is the Sigmoid function. Following existing literature, we use inner product (Wang et al., 2021; Zheng et al., 2021) as the default DECODER.
@@
2.3 The Cold-Start Problem
The cold-start problem is prevalent in various domains and scenarios. In recommendation systems (Chen et al., 2020; Lu et al., 2020; Hao et al., 2021; Zhu et al., 2019; Volkovs et al., 2017; Liu & Zheng, 2020), cold-start refers to the lack of sufficient interaction history for new users or items, which makes it challenging to provide accurate recommendations. Similarly, in the context of GNNs, the cold-start problem refers to performance in tasks involving cold nodes, which have few or no neighbors in the graph. As illustrated in  Figure 1, GNNs usually struggle with cold nodes in LP tasks due to unreliable or missing neighbors’ information. In this work, we focus on enhancing LP performance for cold nodes, specifically predicting the presence of links between a cold node $v\in\mathcal{V}_{cold}$ and target node $u\in\mathcal{V}$ (w.l.o.g.). Additionally, we aim to maintain satisfactory LP performance for warm nodes. Prior studies on cold-start problems (Tang et al., 2020b; Liu et al., 2021; Zheng et al., 2021) inspired this research direction.
@@ 3 Node Duplication to Improve Cold-start Performance
We make a simple but powerful observation: cold nodes are strongly under-represented in the LP training. Given that they have few or even no directly connected neighbors, they hardly participate in the standard supervised LP training as described in Section 2. For example, a model will not see an isolated node unless it is randomly sampled as a negative training edge for another node. In light of such observations, our proposed augmentation technique is simple: we duplicate under-represented cold nodes. By both training and aggregating with the edges connecting the cold nodes with their duplications, cold nodes are able to gain better visibility in the training process, which allows the GNN-based LP models to learn better representations. In this section, we introduce NodeDup in detail, followed by comprehensive analyses of why it works from different perspectives.
@@ 3.1 Proposed Method
The implementation of NodeDup can be summarized into four simple steps: (1): duplicate all cold nodes to generate the augmented node set $\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{cold}$, whose node feature matrix is then $\mathbf{X}^{\prime}\in\mathbb{R}^{(N+|\mathcal{V}_{cold}|)\times F}$. (2): for each cold node $v\in\mathcal{V}_{cold}$ and its duplication $v^{\prime}$, add an edge between them and get the augmented edge set =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv^{\prime}}:\forall v\in\mathcal{V}_{cold}\}$. (3): include the augmented edges into the training set and get $\mathcal{Y}^{\prime}=\mathcal{Y}\cup\{y_{vv^{\prime}}=1:\forall v\in\mathcal{V}_{cold}\}$. (4): proceed with the standard supervised LP training on the augmented graph $G^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime},\textbf{X}^{\prime}\}$ with augmented training set $\mathcal{Y}^{\prime}$. We also summarize this whole process of NodeDup in Algorithm 1 in Appendix C.
@@
Time Complexity. We discuss complexity of our method in terms of the training process on the augmented graph. We use GSage (Hamilton et al., 2017) and inner product decoder as the default architecture when demonstrating the following complexity (w.l.o.g). With the augmented graph, GSage has a complexity of $O(R^{L}(N+|\mathcal{V}_{cold}|)D^{2})$, where $R$ represents the number of sampled neighbors for each node, $L$ is the number of GSage layers (Wu et al., 2020), and $D$ denotes the size of node representations. In comparison to the non-augmented graph, NodeDup introduces an extra time complexity of $O(R^{L}|\mathcal{V}_{cold}|D^{2})$. For the inner product decoder, we incorporate additionally $|\mathcal{V}_{cold}|$ positive edges and also sample $|\mathcal{V}_{cold}|$ negative edges into the training process, resulting in the extra time complexity of the decoder as $O((M+|\mathcal{V}_{cold}|)D)$. Given that all cold nodes have few ($R\leq 2$ in our experiments) neighbors, and GSage is also always shallow (so $L$ is small) (Zhao & Akoglu, 2019), the overall extra complexity introduced by NodeDup is $O(|\mathcal{V}_{cold}|D^{2}+|\mathcal{V}_{cold}|D)$.

### v2

@@ Abstract
Graph Neural Networks (GNNs) are prominent in graph machine learning and have shown state-of-the-art performance in Link Prediction (LP) tasks. Nonetheless, recent studies show that GNNs struggle to produce good results on low-degree nodes despite their overall strong performance. In practical applications of LP, like recommendation systems, improving performance on low-degree nodes is critical, as it amounts to tackling the cold-start problem of improving the experiences of users with few observed interactions. In this paper, we investigate improving GNNs’ LP performance on low-degree nodes while preserving their performance on high-degree nodes and propose a simple yet surprisingly effective augmentation technique called NodeDup. Specifically, NodeDup duplicates low-degree nodes and creates links between nodes and their own duplicates before following the standard supervised LP training scheme. By leveraging a “multi-view” perspective for low-degree nodes, NodeDup shows significant LP performance improvements on low-degree nodes without compromising any performance on high-degree nodes. Additionally, as a plug-and-play augmentation module, NodeDup can be easily applied on existing GNNs with very light computational cost. Extensive experiments show that NodeDup achieves 38.49%, 13.34%, and 6.76% relative improvements on isolated, low-degree, and warm nodes, respectively, on average across all datasets compared to GNNs and the existing cold-start methods.
@@ 1 Introduction
Link prediction (LP) is a fundamental task of graph-structured data (Liben-Nowell & Kleinberg, 2007; Trouillon et al., 2016), which aims to predict the likelihood of the links existing between two nodes in the network. It has wide-ranging real-world applications across different domains, such as friend recommendations in social media (Sankar et al., 2021; Tang et al., 2022; Fan et al., 2022), product recommendations in e-commerce platforms (Ying et al., 2018; He et al., 2020), knowledge graph completion (Li et al., 2023; Vashishth et al., 2020; Zhang et al., 2020), and chemical interaction prediction (Stanfield et al., 2017; Kovács et al., 2019; Yang et al., 2021).
@@
In recent years, graph neural networks (GNNs) (Kipf & Welling, 2016a; Veličković et al., 2017; Hamilton et al., 2017) have been widely applied to LP, and a series of cutting-edge models have been proposed (Zhang & Chen, 2018; Zhang et al., 2021; Zhu et al., 2021; Zhao et al., 2022b). Most GNNs follow a message-passing scheme (Gilmer et al., 2017) in which information is iteratively aggregated from neighbors and used to update node representations accordingly. Consequently, the success of GNNs usually heavily relies on having sufficient high-quality neighbors for each node (Zheng et al., 2021; Liu et al., 2021). However, real-world graphs often exhibit long-tailed distribution in terms of node degrees, where a significant fraction of nodes have very few neighbors (Tang et al., 2020b; Ding et al., 2021; Hao et al., 2021). For example, Figure˜1 shows the long-tailed degree distribution of the Citeseer dataset. Moreover, LP performances w.r.t. node degrees on this dataset also clearly indicate that GNNs struggle to generate satisfactory results for nodes with low or zero degrees. For simplicity, in this paper, we refer to the nodes with low or zero degrees as cold nodes and the nodes with higher degrees as warm nodes.
To boost GNNs’ performance on cold nodes, recent studies have proposed various training strategies (Liu et al., 2020; 2021; Zheng et al., 2021; Hu et al., 2022) and augmentation strategies (Hu et al., 2022; Rong et al., 2019; Zhao et al., 2022b) to improve representation learning quality. For instance, ColdBrew (Zheng et al., 2021) posits that training a powerful MLP can rediscover missing neighbor information for cold nodes; TailGNN (Liu et al., 2021) utilizes a cold-node-specific module to accomplish the same objective. However, such advanced training strategies (e.g., ColdBrew and TailGNN) share a notable drawback: they are trained with a bias towards cold nodes, which then sacrifices performance on warm nodes (empirically validated in Table˜1). However, in real-world applications, both cold nodes and warm nodes are critical (Clauset et al., 2009). On the other hand, while augmentation methods such as LAGNN (Liu et al., 2022b) do not have such bias, they primarily focus on improving the overall performance of GNNs in LP tasks, which may be dominated by warm nodes due to their higher connectivity. Additionally, the augmentation methods usually introduce a significant amount of extra computational costs (empirically validated in Figure˜8). In light of the existing work discussed above on improving LP performance for cold nodes, we are naturally motivated to explore the following crucial but rather unexplored research question:
@@
We observe that cold node LP performance usually suffers because they are under-represented in standard supervised LP training due to their few (if any) connections. Given this observation, in this work, we introduce a simple yet effective augmentation method, NodeDup, for improving LP performance on cold nodes. Specifically, NodeDup duplicates cold nodes and establishes edges between each original cold node and its corresponding duplicate. Subsequently, we conduct standard supervised end-to-end training of GNNs on the augmented graph. To better understand why NodeDup is able to improve LP performance for cold nodes, we thoroughly analyze it from multiple perspectives, during which we discover that this simple technique effectively offers a “multi-view” perspective of cold nodes during training. This “multi-view” perspective of the cold nodes acts similarly to an ensemble and drives performance improvements for these nodes. Additionally, our straightforward augmentation method provides valuable supervised training signals for cold nodes and especially isolated nodes. Furthermore, we also introduce NodeDup(L), a lightweight variation of NodeDup that adds only self-loop edges into training edges for cold nodes. NodeDup(L) empirically offers up to a 1.3$\times$ speedup over NodeDup for the training process and achieves significant speedup over existing augmentation baselines. In our experiments, we comprehensively evaluate our method on seven benchmark datasets. Compared to GNNs and the existing cold-start methods, NodeDup achieves 38.49%, 13.34%, and 6.76% relative improvements on isolated, low-degree, and warm nodes, respectively, on average across all datasets. NodeDup also greatly outperforms augmentation baselines on cold nodes, with comparable warm node performance. Finally, as plug-and-play augmentation methods, our methods are versatile and effective with different LP encoders/decoders. They also achieve significant performance in a more realistic inductive setting. Our code can be found at https://github.com/zhichunguo/NodeDup.
@@ 2 Preliminaries
Notation. Let an attributed graph be $G=\{\mathcal{V},\mathcal{E},\textbf{X}\}$, where $\mathcal{V}$ is the set of $N$ nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edges where each $e_{vu}\in\mathcal{E}$ indicates nodes $v$ and $u$ are linked. Let $\textbf{X}\in\mathbb{R}^{N\times F}$ be the node attribute matrix, where $F$ is the attribute dimension. Let $\mathcal{N}_{v}$ be the set of neighbors of node $v$, i.e., $\mathcal{N}_{v}=\{u|e_{vu}\in\mathcal{E}\}$, and the degree of node $v$ is $|\mathcal{N}_{v}|$. We separate the set of nodes $\mathcal{V}$ into three disjoint sets $\mathcal{V}_{iso}$, $\mathcal{V}_{low}$, and $\mathcal{V}_{warm}$ by their degrees based on the threshold hyperparameter $\delta$ [^1] This threshold $\delta$ is set as 2 in our experiments, based on observed performance gaps in LP on various datasets, as shown in Figure 1 and Figure 10. Further reasons for this threshold are detailed in Section C.1.. For each node $v\in\mathcal{V}$, $v\in\mathcal{V}_{iso}$ if $|\mathcal{N}_{v}|=0$; $v\in\mathcal{V}_{low}$ if $0<|\mathcal{N}_{v}|\leq\delta$; $v\in\mathcal{V}_{warm}$ if $|\mathcal{N}_{v}|>\delta$. For ease of notation, we also use $\mathcal{V}_{cold}=\mathcal{V}_{iso}\cup\mathcal{V}_{low}$ to denote the cold nodes, which is the union of Isolated and Low-degree nodes.
LP with GNNs. In this work, we follow the commonly-used encoder-decoder framework for GNN-based LP (Kipf & Welling, 2016b; Berg et al., 2017; Schlichtkrull et al., 2018; Ying et al., 2018; Davidson et al., 2018; Zhu et al., 2021; Yun et al., 2021; Zhao et al., 2022b), where a GNN encoder learns the node representations and the decoder predicts the link existence probabilities given each pair of node representations. Most GNNs follow the message passing design (Gilmer et al., 2017) that iteratively aggregate each node’s neighbors’ information to update its embeddings. Without the loss of generality, for each node $v$, the $l$-th layer of a GNN can be defined as
$\bm{h}_{v}^{(l)}=\text{\small UPDATE}\big{(}\bm{h}_{v}^{(l-1)},\bm{m}_{v}^{(l-1)}\big{)},\text{s.t.}\;\;\bm{m}_{v}^{(l-1)}=\text{AGG}\big{(}\{\bm{h}_{u}^{(l-1)}\}:\forall u\in\mathcal{N}_{v}\big{)},$ (1)
@@
The Cold-Start Problem. The cold-start problem is prevalent in various domains and scenarios. In recommendation systems (Chen et al., 2020; Lu et al., 2020; Hao et al., 2021; Zhu et al., 2019; Volkovs et al., 2017; Liu & Zheng, 2020), cold-start refers to the lack of sufficient interaction history for new users or items, which makes it challenging to provide accurate recommendations. Similarly, in the context of GNNs, the cold-start problem refers to performance in tasks involving cold nodes, which have few or no neighbors in the graph. As illustrated in  Figure˜1, GNNs usually struggle with cold nodes in LP tasks due to unreliable or missing neighbors’ information. In this work, we focus on enhancing LP performance for cold nodes, specifically predicting the presence of links between a cold node $v\in\mathcal{V}_{cold}$ and target node $u\in\mathcal{V}$ (w.l.o.g.). Additionally, we aim to maintain satisfactory LP performance for warm nodes. Prior studies on cold-start problems (Tang et al., 2020b; Liu et al., 2021; Zheng et al., 2021) inspired this research direction.
@@ 3 Node Duplication to Improve Cold-start Performance
[…]
2: Identify cold node set $\mathcal{V}_{cold}$ based on the node degree.
3:(Step I) Duplicate all cold nodes to generate the augmented node set $\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{cold}$, whose node feature matrix is then $\mathbf{X}^{\prime}\in\mathbb{R}^{(N+|\mathcal{V}_{cold}|)\times F}$.
4:(Step II) Add an edge between each cold node $v\in\mathcal{V}_{cold}$ and its duplication $v^{\prime}$, then get the augmented edge set =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv^{\prime}}:\forall v\in\mathcal{V}_{cold}\}$.
5:(Step III) Add the augmented edges into the training set and get $\mathcal{Y}^{\prime}=\mathcal{Y}\cup\{y_{vv^{\prime}}=1:\forall v\in\mathcal{V}_{cold}\}$.
6:(Step IV) # End-to-end supervised training based on the augmented graph $G^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime},\textbf{X}^{\prime}\}$.
7:for $l=1$ to $L$ do
8:  for $v$ in $\mathcal{V^{\prime}}$ do
10:  end for
11:end for
12:for $(i,j)$ in $\mathcal{Y^{\prime}}$ do
14:end for
Algorithm 1 NodeDup.
As described in Section˜2, a model will not see an isolated node unless it is randomly sampled as a negative training edge for another node in standard supervised LP training. In the same vein, all the cold nodes are strongly underrepresented in the LP training, given their few or even no directly connected neighbors. In light of such observations, our proposed augmentation technique is simple: we duplicate under-represented cold nodes. By both training and aggregating with the edges connecting the cold nodes with their duplications, cold nodes are able to gain better visibility in the training process, which allows the GNN-based LP models to learn better representations. In this section, we introduce NodeDup in detail, followed by comprehensive analyses of why it works from different perspectives.
@@ 3.1 Proposed Method
We summarize the entire process of NodeDup in Algorithm˜1, where the key steps are broken down into four simple stages: Step I: Duplicate all cold nodes to generate the augmented node set $\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{cold}$, whose node feature matrix is then $\mathbf{X}^{\prime}\in\mathbb{R}^{(N+|\mathcal{V}_{cold}|)\times F}$; Step II: For each cold node $v\in\mathcal{V}_{cold}$ and its duplication $v^{\prime}$, add an edge between them and get the augmented edge set =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv^{\prime}}:\forall v\in\mathcal{V}_{cold}\}$; Step III: Include the augmented edges into the training set and get $\mathcal{Y}^{\prime}=\mathcal{Y}\cup\{y_{vv^{\prime}}=1:\forall v\in\mathcal{V}_{cold}\}$; Step IV: Proceed with the standard supervised LP training on the augmented graph $G^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime},\textbf{X}^{\prime}\}$ with augmented training set $\mathcal{Y}^{\prime}$. Based on extensive experimental analysis, we choose to duplicate code nodes once. The impact of both the type of duplicated nodes and the duplication frequency is further analyzed in Section˜4.7.
@@
Time Complexity. We discuss complexity of our method in terms of the training process on the augmented graph. We use GSage (Hamilton et al., 2017) and inner product decoder as the default architecture when demonstrating the following complexity (w.l.o.g). With the augmented graph, GSage has a complexity of $O(R^{L}(N+|\mathcal{V}_{cold}|)D^{2})$, where $R$ represents the number of sampled neighbors for each node, $L$ is the number of GSage layers (Wu et al., 2020), and $D$ denotes the size of node representations. In comparison to the non-augmented graph, NodeDup introduces an extra time complexity of $O(R^{L}|\mathcal{V}_{cold}|D^{2})$. For the inner product decoder, we incorporate additionally $|\mathcal{V}_{cold}|$ positive edges and also sample $|\mathcal{V}_{cold}|$ negative edges into the training process, resulting in the extra time complexity of the decoder as $O(|\mathcal{V}_{cold}|D)$. Given that all cold nodes have few ($R\leq 2$ in our experiments) neighbors, and GSage is also always shallow (so $L$ is small) (Zhao & Akoglu, 2019), the overall extra complexity introduced by NodeDup is $O(|\mathcal{V}_{cold}|D^{2}+|\mathcal{V}_{cold}|D)$.
3.2 How does Node Duplication Help Cold-start LP?
In this subsection, we analyze how such a simple method can improve cold-start LP from two perspectives: the neighborhood aggregation in GNNs and the supervision signal during training, in comparison to self-loops in GNNs (e.g., the additional self-connection in the normalized adjacency matrix by GCN).
Figure 2: Comparison of aggregation mechanisms: GNN with Self-loop, NodeDup for Low-degree nodes, and NodeDup for Isolated nodes.
Aggregation. As described in Equation˜1, when UPDATE()($\cdot$)( ⋅ ) and AGG()($\cdot$)( ⋅ ) do not share the transformation for node features, GNNs would have separate weights for self-representation and neighbor representations, as shown in Figure˜2. The separate weights enable the neighbors and the node itself to play distinct roles in the UPDATE step. By leveraging this property, with NodeDup, the model can leverage the two “views” for each node: first, the existing view is when a node is regarded as the anchor node during message passing, and the additional view is when that node is regarded as one of its neighbors thanks to the duplicated node from NodeDup. Taking the official PyG (Fey & Lenssen, 2019) implementation of GSage (Hamilton et al., 2017) as an example, it updates node representations using $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}+\text{W}_{2}\bm{m}_{v}^{(l)}$. Here, $\text{W}_{1}$ and $\text{W}_{2}$ correspond to the self-representation and neighbors’ representations, respectively. Without NodeDup, isolated nodes $\mathcal{V}_{iso}$ have no neighbors. Thus, the representations of all $v\in\mathcal{V}_{iso}$ are only updated by $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}$. With Step II in NodeDup, the updating process for isolated node $v$ becomes $\bm{h}_{v}^{(l+1)}=\text{W}_{1}\bm{h}_{v}^{(l)}+\text{W}_{2}\bm{h}_{v}^{(l)}=(\text{W}_{1}+\text{W}_{2})\bm{h}_{v}^{(l)}$. It indicates that $\text{W}_{2}$ is also incorporated into the node updating process for isolated nodes, which offers an additional perspective for isolated nodes’ representation learning. Similarly, GAT (Veličković et al., 2017) updates node representations with $\bm{h}_{v}^{(l+1)}=\alpha_{vv}\Theta\bm{h}^{(l)}_{v}+\sum_{u\in\mathcal{N}_{v}}\alpha_{vu}\Theta\bm{h}^{(l)}_{u}$, where $\alpha_{vu}=\frac{\text{exp}(\text{LeakyReLU}(\bm{a}^{\top}[\Theta\bm{h}^{(l)}_{v}||\Theta\bm{h}^{(l)}_{u}]))}{\sum_{i\in\mathcal{N}_{v}\cup v}\text{exp}(\text{LeakyReLU}(\bm{a}^{\top}[\Theta\bm{h}^{(l)}_{v}||\Theta\bm{h}^{(l)}_{i}]))}$. Attention scores in $\bm{a}$ partially correspond to the self-representation $\bm{h}_{v}$ and partially to neighbors’ representation $\bm{h}_{u}$. In this case, neighbor information offers a different perspective compared to self-representation. Such “multi-view” enriches the representations learned for the isolated nodes in a similar way to how ensemble methods work (Allen-Zhu & Li, 2020). Apart from addressing isolated nodes, the same mechanism and multi-view perspective also apply to Low-degree nodes.
Figure 3: Comparison of supervision mechanisms.
Supervision. In LP tasks, edges not only facilitate aggregation but also act as positive supervised training signals, as depicted in Figure˜3. Cold nodes, which typically have few or no positive training edges, are particularly susceptible to out-of-distribution (OOD) issues (Wu et al., 2022), especially in the case of Isolated nodes. Unlike normal self-loops and the self-loops introduced in previous works (Cai et al., 2019; Wang et al., 2020), where self-loops are solely for aggregation, the edges added by NodeDup also serve as positive supervision signals for cold nodes through Step III of Algorithm˜1. By leveraging these additional signals, cold nodes can learn more robust and higher quality embeddings, ultimately improving their performance in LP tasks.